1.  After levying tax on consumers, new supply and demand curve given as Qs=P. Qd=100-(P+T). (Qd stands for quantity demanded, Qs quantity supplies, P is price, T is tax on consumers.

What is unit tax amount T when maximizing government tax revenue?  

 


1
Expert's answer
2021-10-26T09:34:14-0400

When maximizing government tax revenue, the demand and supply should be at equilibrium


At equilibrium quantity supplied equal quantity demanded "Q_s = Q_d"


Therefore:


"P = 100 -(P + T)"


"P= 100 - P - T"


"T = 100 - 2P"


"\therefore" The unit tax amount is "T = 100 -2P"